# Structure and dynamics of the magnetite(001)/water interface from molecular dynamics simulations based on a neural network potential

### v1

@@ I Introduction
Iron oxide interfaces are common in geochemistry and biology and they play crucial roles in phenomena such as corrosion and nucleation (Frankel et al., 1979-03-30; Björneholm et al., 2016-07-13; Munoz et al., 2015; Weiss and Ranke, 2002; Dictor, 1986; Kirschvink et al., 1992). Various technological areas rely on iron oxide interfaces, particularly in applications related to energy conversion and storage, nanoparticles, spintronics and even biomedicine (Tartaj et al., 2011; Sivula et al., 2011; Wu et al., 2015-04-28; Gupta and Gupta, 2005-06; Yanase and Siratori, 1984). The importance of these interfaces and the need for a deeper understanding have driven much experimental and computational research on this topic (see (Parkinson, 2016) for an extensive review on iron oxides). The interface between magnetite ($\rm Fe_{3}O_{4}$) and water is a prominent example of iron oxide interfaces (Zaki et al., 2018; Kraushofer et al., 2019-10-21; Ryan et al., 2022; Liu et al., 2021; Liu and Valentin, 2018-06-14; Weiss and Ranke, 2002; Li and Paier, 2016-01-21; Omranpour et al., 2024). The bare (001) and (111) surfaces of magnetite have been found to be the most stable (Kovács and Lo, 2010; Santos-Carballal et al., 2014) and, more recently, Bliem et al. (Bliem et al., 2014) discovered that the (001) surface undergoes a reconstruction known as the Subsurface Cation Vacancy (SCV) model (see Fig.2), which involves the three outermost layers of magnetite. Later, water agglomeration on the SCV has been studied experimentally and computationally via Density Functional Theory (DFT) by Meier et al. (Meier et al., 2018). Their study showed that energetically favorable configurations for various water coverages on the SCV (up to 8 $\rm H_{2}O$/u.c.) are stabilized by the presence of a partially dissociated water. We label as “dissociated” the OH group formed by dissociated water absorbed on a lattice oxygen and we refer to the “dissociation site” the active site where it takes place (se Fig. 2).
@@ II.3 Simulations details and analysis
Figure 2: Top view of the SCV model. The surface unit cell ($\sqrt{2}\times\sqrt{2}$ rotated $45^{\circ}$ with respect to the bulk) is enclosed in the square and two times replicated. The oxygen atoms are shown in red and the iron atoms in different shades of grey to distinguish between octahedral and tetrahedral irons (light and middle gray) and the subsurface octahedral iron (dark grey). The plus symbols indicate the interstitial tetrahedral irons that during the reconstruction are lifted from the subsurface layer, leaving there the subsurface vacancies (dashed line). The active site (stars) where the water dissociates one hydrogen (“dissociation site”) are located in the tetrahedral iron rows in the site.
@@ III.1 Low water coverage
Experiments including atomic force microscopy and scanning tunneling microscopy in combination with DFT calculations have shown that partially dissociated water clusters of various sizes form at the SCV (Meier et al., 2018). One open question remains about the structure of the water trimer. In experiments, only linear trimers were observed. However, DFT calculations suggested that such linear trimers have the same adsorption energy as triangular trimers. To address this issue, we have simulated the SCV in contact with three water molecules in a unit cell at 300 K. The starting configurations are the two types of trimers on the two exposed SCVs (as shown in Fig. 3) and we ran the system for 54 ns.
These simulations reveal that the linear and triangular trimers exhibit distinct dynamics. To quantify this difference we classified the water structures into clusters of different sizes at each timestep, utilizing a sparse matrix to encode pair distances with a cutoff of 3.5  Å. By identifying clusters of three water molecules (0.98 percent of the time) at the surface and distinguishing between the linear and triangular geometry, we calculated the probability of each trimer conformation. This probability was determined as the ratio of time spent in a particular conformation and the total time spent in any trimer form. Our results averaged from the two surfaces (which agreed up to 0.002) indicate that the linear trimer is 2.55 times more likely to occur than the triangular trimer at 300K:
$\frac{P({\rm linear})}{P({\rm triangular})}=2.55$ (1)
This effect is likely entropic in origin because the potential energy of the two trimer configurations is essentially the same.
@@
Figure 4: Minimum energy configurations of water on the SCV at low coverage. The configurations on the left involve six water molecule per unit cell while on the right eight. The minima in the top row are from (Meier et al., 2018). The other configurations are new stable configurations discovered using the NNP. The SCV is not shown for a better visualization of the water structure. The iron rows are indicated by the dashed lines and the dissociation site by the stars. The number of dissociated hydrogens in the minima found in  (Meier et al., 2018) (top figures) are two while the new ground states (with $\Delta E\leq 0$) have only one dissociated hydrogen.
@@
Next, we investigated the stable structures of the water on the SCV towards the monolayer. Previous configurations were found  (Meier et al., 2018) to preferentially arrange as shown at the top of Fig. 4 for 6 and 8 water molecules per surface unit cell, respectively. We ran MD simulations at 200 K and 300 K for systems comprising one unit cell with 6 and 8 $\rm H_{2}O$ starting from both random configurations and the stable structure reported in Ref. (Meier et al., 2018). During the MD runs, the systems with random initial structures relaxed and one water molecule dissociated while the systems initiated with the stable configurations eventually drifted away from the initial state. After gathering a total of 1522 ns from five independent trajectories for each of the two temperatures and coverages, we picked configurations equally spaced in time. Since two surfaces are involved in the simulation (one on either side of the magnetite slab), two independent water arrangements were produced for each configuration. Then, the obtained structures were relaxed by minimizing the energy using the NNP. This procedure was computationally very efficient with the NNP, allowing a quick initial energy comparison of a large number of structures (on the order of tens of thousands). Subsequently, the 100 lowest energy structures were minimized via DFT. All the candidates which showed lower energy than the already established configurations were selected and their water structures were sliced out and placed on top of the exact same magnetite slab in order to isolate the water contribution. Finally, these setups were further minimized via DFT. Following this procedure, we found new configurations, shown in the middle and bottom row in Fig. 4 for 6 $\rm H_{2}O$/u.c. (left) and for 8 $\rm H_{2}O$/u.c. (right), with lower ($\Delta E=-0.01$ eV and $\Delta E=-0.04$ eV respectively) and comparable energy ($\Delta E=0$) to the previously reported ones (top). Surprisingly, although almost energetically equivalent, the new structures are very different. While the previous minima have two dissociations and two water ions per unit cell, the new ground states have only one dissociated water molecule and one water ion per unit cell.
@@
Figure 6: Mean square displacement MSD (total, parallel and orthogonal to the iron rows) of water oxygens (full line) and hydrogens (dashed llines) at the magnetite SCV interface with 6 water molecules per unit cell. The results are shown for the two surfaces of the slab.

### v2

@@ I Introduction
Iron oxide interfaces are common in geochemistry and biology and they play crucial roles in phenomena such as corrosion and nucleation (Frankel et al., 1979-03-30; Björneholm et al., 2016-07-13; Munoz et al., 2015; Weiss and Ranke, 2002; Dictor, 1986; Kirschvink et al., 1992). Various technological areas rely on iron oxide interfaces, particularly in applications related to energy conversion and storage, nanoparticles, spintronics and even biomedicine (Tartaj et al., 2011; Sivula et al., 2011; Wu et al., 2015-04-28; Gupta and Gupta, 2005-06; Yanase and Siratori, 1984). The importance of these interfaces and the need for a deeper understanding have driven much experimental and computational research on this topic (see (Parkinson, 2016) for an extensive review on iron oxides). The interface between magnetite ($\rm Fe_{3}O_{4}$) and water is a prominent example of iron oxide interfaces (Zaki et al., 2018; Kraushofer et al., 2019-10-21; Ryan et al., 2022; Liu et al., 2021; Liu and Valentin, 2018-06-14; Weiss and Ranke, 2002; Li and Paier, 2016-01-21; Omranpour et al., 2024). The bare (001) and (111) surfaces of magnetite have been found to be the most stable (Kovács and Lo, 2010; Santos-Carballal et al., 2014) and, more recently, Bliem et al. (Bliem et al., 2014) discovered that the (001) surface undergoes a reconstruction known as the Subsurface Cation Vacancy (SCV) model (see Fig.2), which involves the three outermost layers of magnetite. Later, water agglomeration on the SCV has been studied experimentally and computationally via Density Functional Theory (DFT) by Meier et al. (Meier et al., 2018). Their study showed that energetically favorable configurations for various water coverages on the SCV (up to 8 $\rm H_{2}O$/u.c.) are stabilized by the presence of a partially dissociated water. We label as “dissociated” the OH group formed by dissociated water absorbed on a lattice oxygen and we refer to the “dissociation site” the active site where it takes place (see Fig. 2).
@@ II.3 Simulations details and analysis
Figure 2: Top view of the SCV model. The surface unit cell ($\sqrt{2}\times\sqrt{2}$ rotated $45^{\circ}$ with respect to the bulk) is enclosed in the square and two times replicated. The oxygen atoms are shown in red and the iron atoms in different shades of grey to distinguish between octahedral and tetrahedral irons (light and middle gray) and the subsurface octahedral irons (dark grey). The plus symbols indicate the interstitial tetrahedral irons that during the reconstruction are lifted from the subsurface layer, leaving there the subsurface vacancies (dashed circles). The active site (stars) where the water dissociates one hydrogen (“dissociation site”) are located in the tetrahedral iron rows in the site free of irons.
@@ III.1 Low water coverage
Experiments including atomic force microscopy and scanning tunneling microscopy in combination with DFT calculations have shown that partially dissociated water clusters of various sizes form at the SCV (Meier et al., 2018). One open question remains about the structure of the water trimer. In experiments, only linear trimers were observed. However, DFT calculations suggested that such linear trimers have the same adsorption energy as triangular trimers. To address this issue, we have simulated the SCV in contact with three water molecules in one unit cell and in a replicated two times two unit cell, both at constant 300 K. The starting water configurations are the two types of trimers on the two exposed SCVs (as shown in Fig. 3). The systems ran 54 ns and 33.5 ns respectively for the unit cell and the replicated one.
These simulations reveal that the linear and triangular trimers exhibit distinct dynamics. To quantify this difference we classified the water structures into clusters of different sizes at each timestep, utilizing a sparse matrix to encode pair distances with a cutoff of 3.5  Å. By identifying clusters of three water molecules at the surface and distinguishing between the linear and triangular geometry, we calculated the probability of each trimer conformation. This probability was determined as the ratio of time spent in a particular conformation and the total time spent in any trimer form. While the results for the two independent surfaces agreed up to 0.01, for the two systems are very different from each other: the probability of being in a linear trimer is 0.72 for the three water in one unit cell and 0.97 for the replicated system with three water. This discrepancy indicates a large finite size effect due the interaction of the water cluster with its periodic image. Note that the water coverage of the two system also is also different (3 $\rm H_{2}O/u.c$ and 0.75 $\rm H_{2}O/u.c$ ). If we consider isolated trimers at the surface then we can write that:
$\frac{P({\rm linear})}{P({\rm triangular})}\sim 32$ (1)
that is the linear trimer is more then thirty times more likely to occur than the triangular trimer at 300K. This effect is likely entropic in origin because the potential energy of the two trimer configurations is essentially the same.
@@
Figure 4: Minimum energy configurations of water on the SCV at low coverage. The configurations on the left involve six water molecule per unit cell while on the right eight. The minima in the top row are from (Meier et al., 2018) the ones in the bottom row are new stable configurations discovered using the NNP. The SCV is not shown for a better visualization of the water structure but the iron rows are indicated by the dashed lines and the dissociation site by the stars. The number of dissociated hydrogens in the minima found in are two while the new ground states (with $\Delta E<0$) have only one dissociated hydrogen.
@@
Next, we investigated the stable structures of the water on the SCV towards the monolayer. Previous configurations were found  (Meier et al., 2018) to preferentially arrange as shown at the top of Fig. 4 for 6 and 8 water molecules per surface unit cell, respectively. We ran MD simulations at 200 K and 300 K for systems comprising one unit cell with 6 and 8 $\rm H_{2}O$ starting from both random configurations and the stable structure reported in Ref. (Meier et al., 2018). During the MD runs, the systems with random initial structures relaxed and one water molecule dissociated while the systems initiated with the stable configurations eventually drifted away from the initial state. After gathering a total of 1522 ns from five independent trajectories for each of the two temperatures and coverages, we picked configurations equally spaced in time. Since two surfaces are involved in the simulation (one on either side of the magnetite slab), two independent water arrangements were produced for each configuration. Then, the obtained structures were relaxed by minimizing the energy using the NNP. This procedure was computationally very efficient with the NNP, allowing a quick initial energy comparison of a large number of structures (on the order of tens of thousands). Subsequently, the 100 lowest energy structures were minimized via DFT. All the candidates which showed lower energy than the already established configurations were selected and their water structures were sliced out and placed on top of the exact same magnetite slab in order to isolate the water contribution. Finally, these setups were further minimized via DFT. Following this procedure, we found new configurations, shown in the bottom row in Fig. 4 for 6 $\rm H_{2}O$/u.c. (left) and for 8 $\rm H_{2}O$/u.c. (right), with comparable and lower energies ($\Delta E=-0.01$ eV and $\Delta E=-0.9$ eV respectively) to the previously reported ones (top row). The new structures are very different and while the previous minima have two dissociations and two water ions per unit cell, the new ground states have only one dissociated water molecule and one water ion per unit cell. The adsorption energy is defined as
$E_{\text{ads}}(n)=E_{\text{tot}}-E_{\text{slab}}-nE_{H_{2}O},$
where $n$ is the number of water molecules in the system, and it contains the bond energy of the water-magnetite and water-water interactions. The adsorption energy was computed for the new minima:
$E_{\text{ads}}(6)=-6.21\text{ eV},\quad E_{\text{ads}}(8)=-7.19\text{ eV}$
@@
Figure 6: Mean square displacement MSD (total, parallel and orthogonal to the iron rows) of water oxygens (full line) and hydrogens (dashed lines) at the magnetite SCV interface with 6 water molecules per unit cell. The results are shown for the two surfaces of the slab.